OpenStudy (anonymous):

The data below shows the average number of text messages sent daily by a group of people: 7, 8, 4, 7, 6, 2, 6, 4, 100, 7, 4, 8, 2, 1, 8 What will be the effect of removing the outlier in the data? a.) The median number of text messages sent daily will increase. b.)The median number of text messages sent daily will decrease. c.)The difference between the maximum and minimum number of text messages sent daily will increase. d.)The difference between the maximum and minimum number of text messages sent daily will decrease.

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (mathmale):

I'd much rather guide you through finding the solution to this problem yourself. Would you mind taking that data and arranging it in ascending order?

OpenStudy (anonymous):

1,2,4,6,7,8,100

OpenStudy (anonymous):

is that it

OpenStudy (anonymous):

Lets set the data in increasing order:

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

1,2,2,4,4,4,6,6,7,7,7,8,8,8,100

OpenStudy (anonymous):

wouldnt it be A.

OpenStudy (anonymous):

The outlier is 100. The median is 6

OpenStudy (anonymous):

Without 100, the median is still 6

OpenStudy (anonymous):

so.. what will be the answer

OpenStudy (mathmale):

@desmonds : Please understand that the OpenStudy Code of Conduct forbits asking for answers. You can request some guidance towards finding the correct answer yourself, but not ask, "so..what will be the answer?" Larryboxaplenty did the right thing in arranging the data in ascending order. Now, would YOU please explain how you find the median of a data set after you've arranged the data in ascending order? If you don't know how to do that, please google "median":.

OpenStudy (anonymous):

The maximum is the highest number. The minimum is the lowest number. The difference between the max and min is the range. With 100: range = 100 - 1 = 99 Without 100: range = 8 - 1 = 7 Without the outlier, the range decreases

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Do you know how to find the median?